# Twists, Humps, and Pebbles: The Uneven Landscape of End-to-end Speech Recognition and Demographics

### v1

@@ Abstract
Current voice recognition approaches use multi-task, multilingual models for speech tasks like Automatic Speech Recognition (ASR) to make them applicable to many languages without substantial changes. However, broad language coverage can still mask performance gaps within languages, for example, across genders. We systematically evaluate multilingual ASR systems on gendered performance gaps. Using two popular models on three datasets in 19 languages across seven language families, we find clear gender disparities. However, the advantaged group varies between languages. While there are no significant differences across groups in phonetic variables (pitch, speaking rate, etc.), probing the model’s internal states reveals a negative correlation between probe performance and the gendered performance gap. I.e., the easier to distinguish speaker gender in a language, the more the models favor female speakers. Our results show that group disparities remain unsolved despite great progress on multi-tasking and multilinguality. We provide first valuable insights for evaluating gender gaps in multilingual ASR systems. We release all code and artifacts at https://github.com/g8a9/multilingual-asr-gender-gap.
Multilingual Speech Models for Automatic Speech Recognition Exhibit Gender Performance Gaps
Giuseppe ${}^{\clubsuit}$, Beatrice ${}^{\spadesuit}$, Dennis ${}^{\diamondsuit}$${}^{\spadesuit}$, Dirk ${}^{\clubsuit}$$^{\clubsuit}$ Bocconi University, Milan, Italy ${}^{\diamondsuit}$ University of Trento, Trento, Italy ${}^{\spadesuit}$ Fondazione Bruno Kessler, Trento, Italy giuseppe.attanasio3@unibocconi.it
@@ 1 Introduction
Recent advancements in Automatic Speech Recognition (ASR) have showcased remarkable capabilities of “one-model-fits-all” multi-task, multilingual solutions (Radford et al., 2022; Communication et al., 2023; Chu et al., 2023). These neural networks can transcribe, translate, or generate speech in an unprecedented number of languages. Consequently, these models have found widespread adoption [^1] Anecdotally, all major cloud providers and countless smaller companies list a tutorial on how to deploy and serve Whisper (Radford et al., 2022) on their platforms. and are being integrated into applications for the general public. [^2] https://techcrunch.com/2023/06/19/gladia-turns-any-audio-into-text-in-near-real-time/
Figure 1: Model error rate gaps between female and male speakers. Average across language families and data sets. Empty shapes show models that advantage women, filled shapes advantage men. Pairs with one shape indicate gap on one group only. Language count per family in brackets. Ideal scores on the “0” line.
However, applicability to many languages should be one of many optimizations. Voice and language production are among the strongest identity traits, widely varying across people, including genders. These differences have physiological and sociocultural roots (e.g., Labov, 1964; Wolfram, 2004; Alim, 2004; Zimman, 2017, 2020). While physiological differences are rooted in biological sex (e.g., vocal tract length), sociocultural differences (e.g., prescribed registers) reflect gender identity. In this paper, we account for both.
Because of their varied features and societal implications, gendered voice and speech differences—usually as binary distinction—have been extensively studied (Coleman, 1976; Busby and Plant, 1995; Hillenbrand and Clark, 2009). However, modeling gender also poses various challenges for ASR, often leading to performance disparities usually favoring male Tatman (2017); Garnerin et al. (2019, 2021) but sometimes female speakers Adda-Decker and Lamel (2005); Feng et al. (2021); Liu et al. (2022). Those performance disparities can lead to unequal service quality, which is especially harmful to already socially disadvantaged groups (Mengesha et al., 2021). However, most of those studies are within one language (usually English). Evaluating the now widely deployed multilingual ASR models frequently overlooks gendered differences, posing fairness concerns.
@@
We are the first to evaluate gendered performance gaps in massively multilingual ASR models systematically. More specifically, we pose two research questions. (Q1) Do multi-task, multilingual ASR models perform equally well across speakers identifying as women, men, or neither of the two? And, if so, (Q2) can we relate performance gasp to (socio)phonetic features or how models encoded gender internally?
To answer Q1, we evaluate two state-of-the-art multilingual ASR models on three datasets (read and spontaneous speech) covering 19 languages from seven language families. Our results show that models do not perform equally well across genders. Whether models favor male or female speakers depends on the dataset and language. Figure 1 shows average error rates across the tested language families. The model performance of speakers who do not identify with either gender is lower than that of those identifying as male. We conduct phonetics and interpretability experiments to understand the reasons for those gaps (Q2). Phonetic analysis shows no significant differences across genders and cannot explain the observed phenomenon. Interpretability through gender probes reveals a negative correlation between female-male gaps and the ability to discern data across genders from a model’s internal representations. This finding suggests that when presented with speech from men and women, models build different internal representations that can serve as a proxy for gender disparities.
@@
We conducted the first extensive evaluation of gender performance differences between two widely used multilingual ASR models and found significant gender gaps. We release code, data, and all artifacts we produce for future research.
@@
When using gender as a variable, we focus on the speaker’s declared identity. We evaluate whether speech from individuals identifying as female, male, or neither of the two are equally processed by multilingual ASR models. Conceptually, this is the ideal outcome, and we measure performance disparities of minority groups (“female”, “other”) versus the socially advantaged group, i.e., men. We define a system to be biased if it risks further contributing allocational (i.e., quality of service, technology less accessible) and representational harms impacting the minority groups, e.g., feeding into stereotypes about inadequacy of women and speech technology or “shrill,” “incorrect” voice (Tallon, 2019). [^3] Voice technology expert Tom Schalk once said that “many issues with women’s voices could be fixed if female drivers were willing to sit through lengthy training… Women could be taught to speak louder and direct their voices towards the microphone.” https://techland.time.com/2011/06/01/its-not-you-its-it-voice-recognition-doesnt-recognize-women/
@@ 2 Background
Multilingual Speech Models.
Contemporary speech recognition systems based on Transformers (Vaswani et al., 2017) outperform the previous transducer-based systems (Li et al., 2022). Among these new systems, a particularly successful subset uses multi-task, multilingual approaches. These architectures process audio and text with separate Transformers and use different communication mechanisms to enable cross-modal interactions. OpenAI’s Whisper (Radford et al., 2022) uses a standard encoder-decoder architecture. Audio and text sequences can be arbitrary—to differentiate among tasks and languages, it uses special decoding tokens. This strategy allows multi-tasking and multilinguality without architectural changes or fine-tuning (see Figure 1, top-right, in Radford et al. (2022)). Similarly, Meta’s SeamlessM4T (Communication et al., 2023) uses a speech encoder for audio and an encoder-decoder transformer for text (Costa-jussà et al., 2022, NLLB).
While these models have shown strong multilingual capabilities, current assessments primarily focus on overall performance metrics but lack, for example, fairness evaluations in speech setups. [^4] Seamless M4T (Communication et al., 2023) has been previously evaluated for gender bias in translation tasks.
@@
Gendered aspects of the voice are one of the most salient individual traits (Kreiman and Sidtis, 2011; Azul, 2015; Zimman, 2021) and have consequently long been studied in linguistics (Zimman, 2020). The anatomy and makeup of the vocal tract do play a role, as it determines pitch range and formant variations, often regarded as the most distinctive phonetic vocal features of (cis)gender men and women. [^5] Cisgender refers to individuals whose gender identity corresponds to the sex assigned at birth (Fuchs and Toda, 2010). However, gender variation in the voice has been shown to arise from several aspects besides physical ones (Oates and Dacakis, 2015; Zimman, 2018; Becker et al., 2022). As studies including transgender individuals show Zimman (2017), sociocultural factors influence vocal use, including which parts of the pitch range are used (Loveday, 1981; Yuasa, 2008), and articulatory practices for sibilant consonants and vowels that are perceived as gender characterizing (Pharao et al., 2014; Podesva et al., 2016; Li, 2017).
For speech technologies, sociodemographic variation has posed challenges to ASR systems (Sawalha and Abu Shariah, 2013; Liu et al., 2022; Rajan et al., 2022; Fucci et al., 2023). Various works have found ASR models for English and French to better recognize male speech and voices (Tatman, 2017; Garnerin et al., 2019, 2021). This effect is often a result of the under-representation of women in the training data (Meyer et al., 2020; Gaido et al., 2020; Garnerin et al., 2020). However, various works have found the reverse to be true. Several studies found comparatively better performance for women in more spontaneous, conversational data in Dutch (Feng et al., 2022) and English (Adda-Decker and Lamel, 2005; Koenecke et al., 2020). They attribute these findings to sociolinguistic factors, like the higher incidence of disfluencies and informal speech in men.
@@ 3 Methodology
This section describes the experimental design (§3) to answer Q1, i.e., whether there are gendered performance differences in multi-task multilingual ASR models. Our results (§4) confirm it.
@@
We experiment with OpenAI’s Whisper (Radford et al., 2022) and Meta’s SeamlessM4T (Communication et al., 2023), two widely used, state-of-the-art multilingual ASR models. See Appendix A.1 for implementation details.
@@
Among other multilingual datasets available for ASR (Gales et al., 2014; Black, 2019; Pratap et al., 2020; Iranzo-Sánchez et al., 2022; Valk and Alumäe, 2021), we use Mozilla Common Voice (Ardila et al., 2020, CV), [^6] We use CV 16.0 from https://huggingface.co/datasets/mozilla-foundation/common_voice_16_0 Fleurs (Conneau et al., 2023) and VoxPopuli (Wang et al., 2021, VoxPopuli). Our choice of datasets is bound to the availability of reliable speakers’ gender information. [^7] Gender is self-declared in CV and Fleurs, and retrieved from https://multimedia.europarl.europa.eu/en in VoxPopuli. Furthermore, these datasets cover various recording and speech conditions: i) read, where speech is typically well-articulated and based on the reading of pre-defined texts (CV and Fleurs), and ii) spontaneous conditions elicited from public talks and speeches, which allow for more speaker-dependent articulations and variation in word usage (Gabler et al., 2023, VoxPopuli). The type of recording conditions might influence the size of performance disparities, with spontaneous speech providing a more realistic image of the impact of potential disparities for real-use ASR cases.
To avoid unreliable results due to data contamination, we base our analysis on the concatenation of validation and test splits of each dataset. See Appendix A.4 for a discussion on data contamination and how transcription performance varies across splits in our setup.
@@
We include 19 languages from CV and Fleurs, and a subset of those—11 in total, due to data availability—from VoxPopuli. These languages represent diverse language families and data availability conditions (i.e., high-low resource). Our language selection depended on sufficient utterances stratified across genders to allow for proper comparisons. See Table 4-6 for full statistics on datasets in Appendix A.5.
@@ 3.1 Evaluation
We use standard Word- and Character-Error Rate (WER and CER, respectively) to evaluate transcription quality. Following Radford et al. (2022), we i) report WER for all languages but Yoruba and Japanese—where variability in orthography and ambiguous word units may affect evaluation Rowlands (1954); Matsuoka et al. (1997)—for which we use CER, ii) normalize all references and hypotheses texts by transliterating them into Cyrillic for Russian and Serbian, and iii) when evaluating Whisper, we apply the official normalization routines to texts. [^8] Found at https://github.com/openai/whisper/tree/main/whisper/normalizers Moreover, we trim audio to the initial 30 seconds  [^9] Most of the recordings in our datasets are shorter than 30 seconds. See statistics in Table 4-6 in Appendix A.5. and filter out dataset noise by removing records with an empty reference and silence in the snippet. [^10] When processing Fleurs-es, we noticed some records containing only background noise. We used state-of-the-art voice activity detection models to detect and filter them out. See Appendix A.2 for details.
When evaluated in terms of overall quality, both Whisper and Seamless showcase competitive results, respectively, on: Fleurs (12.68 and 19.87 – avg. 19 langs); CV (17.51 and 16.42 – avg. 19 langs); VP (13.59 and 12.74 – avg. 11 langs). For disaggregated WER scores per each language and dataset, see Figure 7 in Appendix A.1.
@@
We use a bootstrapped approach (Søgaard et al., 2014) to estimate $\phi(\cdot)$. We run $n=1000$ iterations sampling 40% of the set. When sampling, we balance the two groups in several records for a reliable comparison. We compute $E(\cdot)$ on the arithmetic means of $\phi(\cdot)$ across the $n$ runs and use a two-sided Student’s t-test to compute the statistical difference between the means.
Filtering and bootstrap sampling improve the evaluation robustness of the skewed speaker and gender distributions. See Appendix A.4 for more details on comparing sampling performance vs. complete sets.
Note that sampling requires attributing each record to an individual speaker—information not available in Fleurs. For consistency, we attribute a speaker ID to each record automatically. Specifically, we i) use state-of-the-art speaker verification embedding models to encode each recording, ii) cluster them using HDBSCAN (Campello et al., 2013), and iii) attribute a speaker ID to each cluster. See Appendix A.3 for details on the pipeline.
@@ 4 Gender Performance Gaps
In the following, we report the results of evaluating Whisper and Seamless, focusing on gender performance disparities. We first investigate whether models are equally able to recognize speech for the female and male group (§4.1), then for speakers that identify as neither of the two comparatively to the male gender group (§4.2).
@@
[…]
@@
Figure 2 reports $E(r_{F},r_{M})$, i.e., the gender gap results for Whisper and Seamless on each dataset. Concerning our Q1, a bird’s eye view immediately attests that multilingual models do not perform equally across female and male speakers, showing a preference for one gender over the other for most instances. Our analysis does not consistently reveal models disadvantaging the feminine group—see values below 0. This finding stands in stark contrast with mounting evidence of a strong masculine bias affecting a wide range of Natural Language Processing (NLP) tasks (e.g., Sun et al., 2019; Stanczak et al., 2022). It is especially noteworthy given the well-known under-representation of women in current resources used for model training (Garnerin et al., 2019; Zanon Boito et al., 2022; Sun and Peng, 2021). We observe that both speech models—albeit to varying degrees—better recognize female speech on average in two out of the three considered datasets (i.e., Fleurs and VP).
Upon closer examination, we also observe quite different behaviors across datasets. First, despite the higher degree of spoken language variation to be expected for the spontaneous VP dataset (Figure 1(a)), the evaluation of this elicited resource shows a comparatively reduced gender gap, with overlapping values between Whisper and Seamless for most languages (i.e., hu, es, fr, nl, cs, de). Conversely, it is on read resource that we attest to a higher degree of variability across models as well as languages, with occasional wide performance disparities toward either females or males in CV (1(c)). Notably, Seamless remains closer to performance equality for most instances on Fleurs (1(b)), i.e., values between 11 and -10.
As such, in line with previous research (see §2), analysis of gender disparities for the ASR reveals a complex picture, which we further analyze and unpack in the following sections.

### v2

# Twists, Humps, and Pebbles: Multilingual Speech Recognition Models Exhibit Gender Performance Gaps

@@ Abstract
Current automatic speech recognition (ASR) models are designed to be used across many languages and tasks without substantial changes. However, this broad language coverage hides performance gaps within languages, for example, across genders. Our study systematically evaluates the performance of two widely used multilingual ASR models on three datasets, encompassing 19 languages from eight language families and two speaking conditions. Our findings reveal clear gender disparities, with the advantaged group varying across languages and models. Surprisingly, those gaps are not explained by acoustic or lexical properties. However, probing internal model states reveals a correlation with gendered performance gap. I.e., the easier it is to distinguish speaker gender in a language using probes, the more the gap reduces, favoring female speakers. Our results show that gender disparities persist even in state-of-the-art models. Our findings have implications for the improvement of multilingual ASR systems, underscoring the importance of accessibility to training data and nuanced evaluation to predict and mitigate gender gaps. We release all code and artifacts at https://github.com/g8a9/multilingual-asr-gender-gap.
Twists, Humps, and Pebbles: Multilingual Speech Recognition Models Exhibit Gender Performance Gaps
Giuseppe Attanasio, Beatrice Savoldi, Dennis Fucci, Dirk Hovy  Bocconi University, Milan, Italy  University of Trento, Trento, Italy  Fondazione Bruno Kessler, Trento, Italy giuseppe.attanasio3@unibocconi.it
@@ 1 Introduction
Figure 1: Study overview. We transcribe three speech datasets across 19 languages and compute gender performance gaps. Next, investigate the models for possible causes of these gaps.
A new class of multi-task, multilingual neural networks (Radford et al., 2022; Communication et al., 2023; Chu et al., 2023) has recently pushed the boundaries of several speech-related tasks, including automatic speech recognition (ASR). As these models offer support to an increasing number of languages at no cost, they have found widespread adoption and have been integrated into applications for the general public, such as real-time voice transcription. [^1] E.g., http://tcrn.ch/4el8Brr However, while the user base expands, one question has yet to be answered: will usability be the same for everyone? In this paper, we seek an answer to this question by studying how systems understand the voices of different genders and across multiple languages. These two axes of analysis are motivated by complementary interests.
On the one hand, voice and language production are among the strongest identity traits. They vary across individuals, sociodemographic groups (Labov, 1964; Wolfram, 2004; Alim, 2004), and, crucially, across genders. Gendered differences in the voice—rooted in physiological factors linked to biological sex (e.g., vocal tract length), as well as in sociocultural ones (e.g., prescribed registers) (Zimman, 2017, 2020)—have been extensively studied in sociolinguistics (e.g., Coleman, 1976; Busby and Plant, 1995; Hillenbrand and Clark, 2009). Also, empirical studies have shown that gender affects the performance of traditional ASR systems in English (Garnerin et al., 2019). On the other hand, voice and speech can change drastically across languages, cultures, and in function of gender. For example, males’ and females’ pitch—a measure of the highness of the voice—is much closer in Japanese- than in English-speaking people (Loveday, 1981; Yuasa, 2008).
Despite a large literature on gender, voice, and ASR, to our knowledge, no studies have tested whether multi-task, multilingual ASR models serve genders to a comparable level across languages. Arguably, overlooking potential disparities can result in unequal service quality for already socially disadvantaged individuals (Mengesha et al., 2021; Tatman, 2017).
@@
To fill this gap, we systematically study gendered performance gaps in massively multilingual ASR models. We set out to answer two research questions. (Q1) Do multi-task, multilingual ASR models perform equally well across speakers identifying as women, men, or neither of the two? If so, (Q2), can we relate gender gaps to acoustic and lexical variation in data or model internal states?
To answer Q1, we evaluate two state-of-the-art multilingual open-weight ASR models on three datasets, covering 19 languages and two speech conditions (i.e. read and spontaneous). The results of our extensive evaluation show that models systematically exhibit gender performance gaps. However, whether models favor male or female speakers depends on the dataset and language. Results on one dataset also highlight subpar performance for speakers who do not identify with either gender compared to males.
While studying acoustic and lexical phenomena in test data (Q2), we found no significant correlations of specific features with performance gaps between male and female speakers. This objective finding underscores the complexity of the issue and the need for further research. Instead, interpretability analyses suggest that when presented with speech from men and women, models build different internal representations that can serve as a proxy for gender disparities.
@@
We conduct the first extensive evaluation of two widely used multilingual ASR models for gender performance differences. We document significant gender gaps by inspecting model internal states. We release code, data, and all artifacts we produce for future research.
@@
When using gender as a variable, we rely on speakers’ declared identity (see §Ethical Considerations). We evaluate whether speech from individuals identifying as female, male, or neither of the two is processed equally well by multilingual ASR models. We measure disparities of minority groups versus the socially advantaged group, i.e., men. Performance parity is the ideal outcome. We define a system to be biased if it risks further contributing allocative (i.e., quality of service, technology less accessible) and representational harms impacting the minority groups, e.g., feeding into stereotypes about the inadequacy of women and speech technology or “shrill,” “incorrect” voice (Tallon, 2019). [^2] As voice technology expert Tom Schalk once put it, “many issues with women’s voices could be fixed if female drivers were willing to sit through lengthy training… Women could be taught to speak louder and direct their voices towards the microphone.” https://bit.ly/time-shalck
@@ 2 Background
Multi-task Multilingual Speech Models.
Contemporary speech recognition systems process audio and text separately and use common cross-modal interactions—e.g., OpenAI’s Whisper (Radford et al., 2022) is loosely inspired to a standard Transformer Vaswani et al. (2017). Crucially, multitasking and multilinguality come at virtually no cost for the user. With Whisper, for example, special “task” and “language” tokens can be prepended to the decoder input to change its functions. This strategy allows multi-tasking and multilinguality without architectural changes or fine-tuning (see Figure 1, top-right, in Radford et al. (2022)). Similarly, Meta’s SeamlessM4T (Communication et al., 2023) uses a speech encoder for audio and an encoder-decoder transformer for text (Costa-jussà et al., 2022, NLLB).
@@
Gendered aspects of the voice are one of the most salient individual traits (Kreiman and Sidtis, 2011; Azul, 2015; Zimman, 2021) and have long been studied in linguistics (Zimman, 2020). The anatomy and makeup of the vocal tract do play a role, as it determines pitch range and formant variations, often regarded as the most distinctive vocal features of cisgender men and women. [^3] Cisgender describes individuals whose gender identity matches their birth-assigned sex (Fuchs and Toda, 2010). However, gender variation in the voice has been shown to arise from several aspects besides physical ones (Oates and Dacakis, 2015; Zimman, 2018; Becker et al., 2022). As studies including transgender individuals show Zimman (2017), sociocultural factors influence vocal use, including which parts of the pitch range are used (Loveday, 1981; Yuasa, 2008), and articulatory practices for sibilant consonants and vowels that are perceived as gender characterizing (Pharao et al., 2014; Podesva et al., 2016; Li, 2017).
For speech technologies, sociodemographic variation has posed challenges to ASR systems (Sawalha and Abu Shariah, 2013; Liu et al., 2022; Rajan et al., 2022; Fucci et al., 2023). Various works have found ASR models for English and French to recognize better male speech and voices (Tatman, 2017; Garnerin et al., 2019, 2021). This effect is often a result of the under-representation of women in the training data (Meyer et al., 2020; Gaido et al., 2020; Garnerin et al., 2020). However, various works have found the reverse to be true. Several studies found comparatively better performance for women in more spontaneous, conversational data in Dutch (Feng et al., 2022) and English (Adda-Decker and Lamel, 2005; Koenecke et al., 2020). They attribute these findings to sociolinguistic factors, like the higher incidence of disfluencies and informal speech in men.
@@ 3 Methodology
This section describes the experimental design to answer Q1, i.e., whether there are gendered performance differences in multi-task multilingual ASR models. Our results (§4) confirm it.
@@
We experiment with OpenAI’s Whisper (Radford et al., 2022, Whisper) and Meta’s SeamlessM4T (Communication et al., 2023, Seamless), two widely used, state-of-the-art multilingual ASR models (details in Appendix A.1).
@@
Among other multilingual datasets available for ASR (Gales et al., 2014; Black, 2019; Pratap et al., 2020; Iranzo-Sánchez et al., 2022; Valk and Alumäe, 2021), we use Mozilla Common Voice (CV, Ardila et al., 2020), [^4] We use CV 16.0 from https://huggingface.co/datasets/mozilla-foundation/common_voice_16_0 Fleurs (Conneau et al., 2023) and VoxPopuli (Wang et al., 2021). Mainly, dataset selection is bound to the availability of reliable speakers’ gender information. We use the gender labels that come with each dataset. [^5] Speakers indicated gender identity with “Male” or “Female” labels in Fleurs and CV. The latter also includes an “Other” label. Gender is retrieved from https://multimedia.europarl.europa.eu/en in VoxPopuli. Moreover, these datasets cover two distinct recording and speech conditions: i) read, where speech is typically well-articulated and based on the reading of pre-defined texts (CV and Fleurs), and ii) spontaneous conditions elicited from public speeches (VoxPopuli), which allow for more speaker-dependent articulations and variation in word usage (Gabler et al., 2023), and thus represent a testbed closer to real-world use cases of ASR technologies. We base our analysis on the concatenation of validation and test splits of each dataset to avoid unreliable results due to training data contamination. See Appendix A.5 for a discussion on data contamination and how transcription performance varies across splits in our setup.
@@
We include 19 languages from CV and Fleurs, and a subset of those—11 in total, due to data availability—from VoxPopuli. They represent eight diverse language families and data availability conditions (i.e., high-low resource). The dataset choice depended on whether sufficient utterances stratified across genders were available for meaningful comparisons. See Tables 5-7, Appendix A.6 for all the statistics.
@@ 3.1 Evaluation
We use standard Word- and Character-Error Rate (WER and CER, respectively) to evaluate transcription quality. Following Radford et al. (2022), we i) report WER for all languages but Yoruba and Japanese—where variability in orthography and ambiguous word units may affect evaluation Rowlands (1954); Matsuoka et al. (1997)—for which we use CER, ii) transliterate all Russian and Serbian references and hypotheses into Cyrillic, and iii) apply the official normalization routines to texts when evaluating Whisper. [^6] Found at https://github.com/openai/whisper/tree/main/whisper/normalizers Moreover, we trim audio to the initial 30 seconds [^7] Most recordings in our datasets are shorter than 30 seconds. See statistics in Table 5-7 in Appendix A.6. and filter out dataset noise by removing records with an empty reference and silence in the snippet. [^8] When processing Fleurs-es, we noticed some records containing only background noise. We used state-of-the-art voice-activity detection models to detect and filter them out. See Appendix A.3 for details.
When evaluated in in terms of of overall quality, both Whisper and Seamless showcase competitive results, respectively, on Fleurs (12.68 and 19.87 – avg. 19 langs); CV (17.51 and 16.42 – avg. 19 langs); VP (13.59 and 12.74 – avg. 11 langs). For disaggregated error rate (ER) scores, see Figure 5 in Appendix A.1.
@@
We use a bootstrapped approach (Koehn, 2004; Søgaard et al., 2014) to estimate $\phi(\cdot)$. We sample $40\%$ of the smallest group and the same number of records from the largest group for $n=1000$ iterations. We sample stratifying on speakers to avoid skewing the speaker distribution. Stratified, gender-balanced sampling ensures a reliable comparison. We compute $E(\cdot)$ on the arithmetic means of $\phi(\cdot)$ across the $n$ runs and use a two-sided Student’s t-test to compute the statistical difference between the means. See Appendix A.5 for more details on comparing sampling performance vs. complete sets.
Note that sampling requires attributing each record to an individual speaker—information not available in Fleurs. For consistency, we attribute a speaker ID to each record automatically. Specifically, we i) use state-of-the-art speaker verification embedding models to encode each recording, ii) cluster them using HDBSCAN (Campello et al., 2013), and iii) assign a speaker ID to each cluster. See Appendix A.4 for details on the pipeline.
@@ 4 Gender Performance Gaps
In the following, we report the results of evaluating Whisper and Seamless, focusing on gender performance disparities. We first investigate whether models are equally able to recognize female and male speakers (§4.1). Then, limited to CV, we repeat the analysis for speakers identifying with neither of the two against males (§4.2).
@@
Figure 2: Model error rate gap (Eq. 1). Positive values indicate better performance on men, negative ones on women. $p<0.05$.
@@
Figure 2 reports $E(r_{F},r_{M})$, i.e., the gender gap results for Whisper and Seamless on each dataset. Concerning our Q1, a broad overview reveals that these models do not perform equally across female and male speakers, often showing a preference for one gender over the other. This gender disparity in multilingual models is a key point of our investigation. Our analysis does not consistently reveal models disadvantaging the feminine group—see values below 0. This finding stands in stark contrast with mounting evidence of a strong masculine bias affecting a wide range of Natural Language Processing (NLP) tasks (e.g., Sun et al., 2019; Stanczak et al., 2022). It is especially noteworthy given the well-known under-representation of women in current resources used for model training (Garnerin et al., 2019; Zanon Boito et al., 2022; Sun and Peng, 2021). We observe that both speech models—albeit to varying degrees—better recognize female speech on average in two out of the three considered datasets (i.e., Fleurs and VP).
Upon closer examination, we also observe substantially different behaviors across datasets. First, despite the higher degree of spoken language variation to be expected in spontaneous recordings from the VP dataset (Figure 2(a)), results shows a comparatively reduced gender gap, with overlapping values between Whisper and Seamless for most languages (i.e., hu, es, fr, nl, cs, de). Conversely, it is on read resource that we attest to a higher degree of variability across models as well as languages, with occasional wide disparities toward either females or males in CV (2(c)). Notably, Seamless remains closer to performance equality for most instances on Fleurs (2(b)), i.e., values between 11 and -10.
As such, in line with previous research (see §2), analysis of (binary) gender disparities for the ASR reveals a complex picture, which we further analyze and unpack in the following sections.
@@
Restricting our analysis to validation and test sets leaves us with few records for speakers who do not identify as either male or female. As such, we report results for five high-resource languages, Catalan, German, English, Spanish, and French. [^10] Size varies from 8 (ca, min) to 86 (es, max) records.
Table 1 reports $E(r_{O},r_{M})$ for both models. Results suggest that models penalize non-male speech and that gaps are larger than disparities observed for the female/male groups. Given the limited sample size, we deem this analysis only exploratory. However, it is a valuable first overview of the model’s behaviors for gender-non-conforming voices. We underscore the need for more representation of diverse voices and sociodemographic groups in current resources, and for which monitoring fairness in existing speech models remains out of reach. See §7 for an expanded discussion.
5 Acoustic and Lexical Analysis
Gendered performance gaps vary across languages and datasets. Here, we dive into a focused analysis on acoustic aspects or lexical phenomena present in the test data (Q2). Due to data availability, we conduct this analysis on the “Female” - “Male” setup.
5.1 Acoustic Analysis